### v1

# Contextual Stochastic Optimization for Order Fulfillment Optimization

[…]
Contextual Stochastic Optimization for Order Fulfillment Optimization
[…]
Contextual Stochastic Optimization for Omnichannel Multi-Courier Order Fulfillment Under Delivery Time Uncertainty
[…]
Tinghan Ye, Sikai Cheng, Amira Hijazi, Pascal Van Hentenryck
[…]
NSF AI Institute for Advances in Optimization (AI4OPT), Georgia Institute of Technology, Atlanta, GA
[…]
Problem definition: The paper studies a large-scale order fulfillment problem for a leading e-commerce company in the United States. The challenge involves selecting fulfillment centers and shipping carriers with observational data only to efficiently process orders from a vast network of physical stores and warehouses. The company’s current practice relies on heuristic rules that choose the cheapest fulfillment and shipping options for each unit, without considering opportunities for batching items or the reliability of carriers in meeting expected delivery dates. Methodology / results: The paper develops a data-driven Contextual Stochastic Optimization (CSO) framework that integrates distributional forecasts of delivery time deviations with stochastic and robust order fulfillment optimization models. The framework optimizes the selection of fulfillment centers and carriers, accounting for item consolidation and delivery time uncertainty. Validated on a real-world data set containing tens of thousands of products, each with hundreds of fulfillment options, the proposed CSO framework significantly enhances the accuracy of meeting customer-expected delivery dates compared to current practices. It provides a flexible balance between reducing fulfillment costs and managing delivery time deviation risks, emphasizing the importance of contextual information and distributional forecasts in order fulfillment. Managerial implications: This is the first paper that studies the omnichannel multi-courier order fulfillment problem with delivery time uncertainty through the lens of contextual optimization, fusing machine learning and optimization. The results provide actionable insights for e-commerce companies and online retailers to enhance service quality and customer satisfaction through efficient order fulfillment strategies that accounts for delivery time uncertainty. Adopting this framework can balance cost effectiveness with customer service, potentially leading to higher customer retention and profitability.
[…]
logistics, order fulfillment, machine learning, contextual stochastic optimization, contextual robust optimization
1 Introduction
Over the past decade, online shopping has surged dramatically. The e-commerce share of total retail sales in the United States has risen from 6% in 2015 to 16% in 2024, exceeding 1.1 trillion dollars (US Department of Commerce 2024). Amid intense competition in pricing, service, and marketing, driven by the lucrative e-commerce market, logistical performance has become a key element to success. Last-mile fulfillment, in particular, is critical as it accounts for 55% of the total transportation costs in the fulfillment process (Viswanathan 2023). This evolving market landscape poses unprecedented challenges for e-commerce companies striving to deliver satisfactory customer services. Fortunately, the growing volume of data also offers opportunities for data-driven operational improvements, enabling companies to boost revenue and cut costs (Fisher et al. 2019).
In addition, after the COVID-19 pandemic, the demand for timely home delivery has also increased significantly. As a result, omnichannel fulfillment and the use of multiple couriers, including crowd-shipping, have become essential components of the fulfillment landscape, drawing significant attention both in industry and academia (Dethlefs et al. 2022, Mohri et al. 2023, Das et al. 2023). The omnichannel approach combines distribution centers (DCs) and brick-and-mortar stores to fulfill online orders. Unlike the traditional method that uses DCs for online orders and stores for in-store purchases separately, the omnichannel approach leverages the proximity of physical stores to delivery addresses. This can potentially offer more cost-effective and faster delivery options for online customers (Acimovic and Farias 2019). Meanwhile, in the crowd-shipping model, retailers recruit non-dedicated drivers, such as gig couriers like Roadie, to handle pick-up and delivery tasks for online customers. This approach can alleviate the workload of employed drivers during peak hours and offer faster delivery to customers, resulting in higher service satisfaction (Zehtabian et al. 2022, Behrendt et al. 2023). The problem considered in this paper involves fulfilling orders by choosing from various types of fulfillment centers, such as warehouses and stores, and different couriers, including traditional shipping companies and gig couriers.
Given that transportation and shipping are significant cost drivers in e-commerce (Dethlefs et al. 2022, Kuhn and Sternbeck 2013, Hübner et al. 2013), delivery consolidation has emerged as an effective strategy for fulfilling orders with multiple items. This approach has been shown to enhance consumer satisfaction (Wagner et al. 2023), while also reducing shipping costs for online retailers (Chen et al. 2024). In addition to cutting parcel expenses, item consolidation enhances the customer experience by eliminating the need to receive multiple packages for a single order. This approach not only improves convenience but also promotes a more sustainable and environmentally friendly fulfillment process by decreasing the number of packages shipped and the number of delivery trips needed. This reduction in shipments and trips leads to lower carbon emissions and less packaging waste, contributing to greener logistics practices (Ülkü 2012). This paper also investigates multi-item order consolidation in its omnichannel fulfillment problem.
Most existing studies on order fulfillment optimization have not explicitly considered the uncertainty in delivery timeliness. However, Freedman (2019) highlight the strong desire of customers for accurate delivery time promises and faster service. The survey indicates that around 44% of online shoppers have abandoned their shopping carts because the items they wanted would not arrive on time, while 20% abandon orders with unclear delivery dates. Furthermore, 28% of shoppers are willing to pay for expensive, expedited shipping to receive products at their desired time. The statistics clearly show that customers prioritize service quality, sometimes even more than the product value, when making purchase decisions. Moreover, Salari et al. (2022) point out the importance of considering delivery time uncertainty in order fulfillment to better meet customer expectations and improve satisfaction. Their proposed promised delivery time policy demonstrates a 6.1% increase in sales volume in numerical experiments. Notably, overpromising delivery times can lead to higher return rates, lower repurchase rates, and dissatisfaction with delayed service. Additionally, customers can have mixed sentiments over underpromised delivery times, as early deliveries may not align with their availability. Therefore, missing the exact promised delivery date can negatively impact short-term and long-term sales, both implicitly and explicitly (Chan et al. 2018, Cui et al. 2020, Salari et al. 2022). Given the crucial role of delivery timeliness, the paper incorporates delivery time uncertainty into the omnichannel multi-courier order fulfillment optimization problem through a contextual stochastic optimization (CSO) framework.
One of the key challenges of Omnichannel Multi-Courier Order Fulfillment Optimization comes from the nature of historical data: only observational data is available, i.e., only the delivery time deviations of the selected options are available in the data set. In addition, in real time, these deviations are only observable once the decisions have been taken. The availability of observation data only raises fundamental challenges that are typically not addressed in prior work.
To address these challenges, this paper proposes a Contextual Stochastic Optimization (CSO) framework, encompassing both risk-neutral and robust approaches. Unlike traditional stochastic programming methods that depend on static, predefined probability distributions for uncertainties, CSO combines machine learning techniques with mathematical optimization to dynamically predict uncertainties based on contextual information. Thanks to the availability of contextual information, CSO has the flexibility of generating more tailored and effective decisions, leveraging the most current and relevant information. Such flexibility is particularly beneficial in the dynamic environment of online retail, where conditions and customer behaviors can change rapidly, requiring a more responsive approach to decision making. The effectiveness of CSO has been validated in the existing operations research and management science literature (Mišić and Perakis 2020, Sadana et al. 2024). This approach has proven successful in various applications, including inventory control (Bertsimas et al. 2016, Meller et al. 2018), price and revenue management (Ito and Fujimaki 2016, Perakis et al. 2023), and, more recently, school redistricting (Guan et al. 2024).
1.1 Contributions
The key novelty in this paper is a CSO framework for an omnichannel multi-courier order fulfillment optimization with observational data only. Existing studies on CSO typically assume the availability of a dataset $\mathcal{D}=\{(\mathbf{s}_{1},\mathbf{c}_{1}),(\mathbf{s}_{2},\mathbf{c}_{2}),\ldots\}$, pairing observed covariates with complete uncertain parameter vectors. However, real-world scenarios often involve high-dimensional uncertain parameters where individual data points capture only partial information, particularly when uncertainty measurement depends on the chosen solution, which is the case in the application considered in this paper.
To address the challenge of observational data, this paper introduces a generalized CSO framework addressing these real-world complexities. The novel CSO framework is applied to a sophisticated stochastic omnichannel order fulfillment optimization problem with multiple carrier options and delivery time uncertainties. In collaborating with a major US online retailer, the framework is rigorously evaluated on an industrial order fulfillment dataset featuring a vast, intricate fulfillment network. This application demonstrates the framework capabilities and its potential impact on large-scale, real-world optimization challenges. The contributions of the paper can be summarized as follows.
  1. 1.
    To the best of the authors’ knowledge, this paper is the first study of the order fulfillment problem in the context of omnichannel and multi-courier logistics through the lens of data-driven contextual stochastic optimization.
  2. 2.
    A mixed-integer linear program (MILP) is developed for a novel order fulfillment optimization problem that jointly considers the locations of fulfillment centers and the selection of shipping carriers for online order fulfillment. This MILP accounts for the consolidation of items within each order.
  3. 3.
    Two classes of machine learning models are proposed to learn the delivery time deviations: probabilistic multi-class classification models and quantile regression models that provide distributional forecasts of delivery time deviations.
  4. 4.
    The proposed MILP and machine learning models are integrated into a generalized CSO framework. The framework integrates the conditional distributions of delivery time deviations with stochastic and robust optimization models via sample average approximation (SAA) and tractable reformulations.
  5. 5.
    The paper applies the CSO framework to a large-scale real-world instance, demonstrating its impact on balancing reduced fulfillment costs and increased on-time order deliveries. The results validate the practical effectiveness of contextual optimization in an e-commerce setting, bringing significant benefits compared to the baselines.
1.2 Structure of the Paper
The remaining sections of the paper are organized as follows. Section 2 reviews the related literature. Section 3 formalizes the generic CSO problem addressed in this paper and proposes SAA and robust optimization methods. The methods are applied to an unique order fulfillment problem in Section 4, which describes the MILP model and the machine learning models. Section 5 describes the case study, provides experimental results, and discusses actionable insights obtained from the experiments based on a real dataset. Section 6 concludes the paper.
2 Related Literature
There are several streams of literature related to this paper in terms of methodology and problem context.
2.1 Data-Driven Decision Making
This paper adds to the growing body of literature on data-driven decision-making, particularly at the intersection of machine learning and optimization, which has led to the development of CSO. Sadana et al. (2024) put forth a comprehensive survey on contextual optimization. Tutorials on various methodologies and applications of CSO models in logistics and operations management are available in the works of Mišić and Perakis (2020), Qi and Shen (2022), and Tian et al. (2023). Among the body of research on CSO, the most relevant to this paper are those that focus on learning conditional distributions of uncertain parameters and incorporating these forecasts into optimization models. Data-driven approaches typically build these forecasts from empirical residual errors (Deng and Sen 2022, Kannan et al. 2023, Perakis et al. 2023) or weighted empirical distributions based on proximity to the training data (Bertsimas and Kallus 2020, Notz and Pibernik 2022). However, these studies assume full observation of uncertain parameters for each data point, which differs from the context of this paper. Recent research has also explored integrating downstream optimization effects into ML model training. Elmachtoub and Grigas (2022) introduce the Smart Predict then Optimize (SPO) loss and its convex surrogate for point prediction models, with subsequent extensions to conditional distributional estimation (Qi et al. 2021, Kallus and Mao 2023). However, the SPO loss framework is not directly applicable to this paper’s context, as the dataset analyzed here is observational. In such datasets, outcomes are only partially observed based on decisions made, presenting unique challenges not addressed by these existing methods.
This paper thus contributes to the field of data-driven decision-making that handles observational data, a domain explored by prior research, e.g., in Bertsimas et al. (2019) and Jo et al. (2021) using optimal prescriptive trees for prescriptive analytics in personalized treatment. While these methods effectively learn policies mapping features to actions, they cannot be directly applied to this fulfillment problem considered in this paper due to the presence of instance-specific hard constraints on decisions. The proposed framework addresses these limitations, offering a novel approach to CSO problems with observational data and constrained decision spaces.
Unlike earlier CSO applications that focus on risk-neutral objectives, a growing stream of research is concerned with risk-averse optimization models that incorporate contextual information (Bayram et al. 2022, Peršak and Anjos 2023, Patel et al. 2024, Sun et al. 2024). Building on these studies, this paper also adapts traditional robust optimization models to tackle a unique CSO problem that differs from those explored in other research.
2.2 Order Fulfillment Problem
The novelty in this paper is the application of CSO to a unique instance of the order fulfillment problem. The order fulfillment problem has been extensively studied, and Acimovic and Farias (2019) provide a tutorial on related algorithms developed for addressing the fulfillment optimization problem. In terms of problem context, the most relevant works are those related to omnichannel and multi-item order fulfillment (Xu et al. 2009, Acimovic and Graves 2015, Jasin and Sinha 2015, Zhao et al. 2020, Wei et al. 2021, Ma 2023, Chen et al. 2024). However, these works do not address fulfillment involving a variety of carrier shipping options.
Among the research on order fulfillment, demand uncertainty is the most commonly studied source of randomness (Jasin and Sinha 2015, Zhao et al. 2020, Das et al. 2023, DeValve et al. 2023, Ma 2023). In this paper, demand uncertainty is not considered because the focus is on a single-period order fulfillment problem, rather than a forward-looking approach that spans multiple periods.
A few studies do address delivery time uncertainty at various stages of the order fulfillment process or throughout the entire process. For instance, Raj et al. (2024) propose an integrated queuing model that accounts for delivery time uncertainty in different parts of the order life cycle. They incorporate these uncertainties into an optimization model that minimizes costs while adhering to a delivery probability constraint to ensure on-time delivery.
In contrast to using queuing models, this study is particularly relevant to data-driven, ML-based methods that address for delivery time uncertainty in order fulfillment. Specifically, Liu et al. (2021) integrate point predictions of travel time with a last-mile delivery order assignment optimization problem. They propose both a SAA model and a distributionally robust optimization (DRO) model, with the latter constructing a moment-based ambiguity set. Similarly, Kandula et al. (2021) explore a last-mile delivery problem, focusing on delivery success as the uncertainty. They map success probabilities to delivery times, which are then integrated into a Vehicle Routing Problem with Time Windows. Bayram et al. (2022) study a robust order batching optimization problem in warehouse picking and packing. They use random forest to predict order processing time to form an uncertainty set within a robust optimization model. This paper builds on similar concepts, developing tractable approximations and reformulations of integer programming-based optimization models. More precisely, the uncertainty in the order fulfillment problem considered in this paper is the potential order delivery deviations, which can occur at various stages of the fulfillment process. A similar uncertainty addressed using data-driven approach can be found in Salari et al. (2022), who develop a quantile regression forest-based method to generate distributional forecasts of order delivery times. Their approach then applies a decision rule, designed to balance the asymmetric costs of early and late deliveries, to produce expected arrival times for customers. This paper differs in two key aspects. First, the distributional forecasts are directly incorporated into contextual optimization models, eliminating the need for a separate decision rule to generate point predictions. Second, the CSO framework proposed in this paper is more comprehensive, accommodating multiple carrier options, whereas Salari et al. (2022) focus on a single-carrier scenario.
3 Problem and Methodology
This section describes a generic contextual stochastic optimization problem (denoted by ${\cal P}$) for fulfillment applications where only an observational dataset is available. It then outlines the overall approach taken to address the problem.
3.1 Problem Description
Problem ${\cal P}$ consists in fulfilling a set of incoming requests sequentially, each of which contains orders for a set of products. Each request $o$ has an associated quantity $q_{io}$ for some product $i\in\mathcal{I}$. These products can be fulfilled from a set $\mathcal{V}$ of selection options. When a request $o$ arrives and $o$ contains a subset of products in $\mathcal{I}$, the goal is to determine the quantity of product $i$ assigned to selection option $v$. The decision variables can thus be captured by a vector $\mathbf{z}_{o}=(z_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, where each $z_{ivo}$ indicates the quantity of product $i$ assigned to selection option $v$ in request $o$. Each such assignment must satisfy operational constraints and the set of feasible assignments is denoted by $\mathcal{Z}_{o}$. For each product $i$, each selection option $v$ is associated with an uncertain cost $\tilde{c}_{ivo}\in\mathbb{R}$. The full vector of uncertain costs, governed by a probability distribution $\mathbb{P}$, is denoted by $\mathbf{\tilde{c}}_{o}=(\tilde{c}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$. The objective function $g(\mathbf{z}_{o},\mathbf{\tilde{c}_{o}})$ of the problem depends on these uncertain costs and the decision variables. Problem ${\cal P}$ is thus a risk-neutral stochastic optimization which can be formulated as follows for request $o$:
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (1)
where $\mathbb{E}_{\mathbf{\tilde{c}}_{o}\sim\mathbb{P}}[\cdot]$ represents the expectation of the objective function over the probability distribution $\mathbb{P}$.
3.2 Contextual Stochastic Optimization (CSO)
In Problem (1), the distribution $\mathbb{P}$ is not readily available. However, side information is typically available for product $i$ and selection option $v$. Let the vector of covariates for request be $\mathbf{s}_{ivo}$. In the context of an omnichannel order fulfillment problem, these covariates may represent the state of the fulfillment network, the weather, and the current demand for products. All these covariate vectors are aggregated into state vector $\mathbf{s}_{o}=(\mathbf{s}_{ivo})_{i\in\mathcal{I},v\in\mathcal{V}}$, which encompasses all contextual information available to the request at the time of the assignment decision.
The distribution $\mathbb{P}$ is then approximated by the conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ of the cost vector given the observed state vector. This leads to the following Contextual Stochastic Optimization (CSO) problem:
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})],$ (2)
where $\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}[\cdot]$ is the conditional expectation of the objective function given the state vector $\mathbf{s}_{o}$. Let the uncertain cost vector and the aggregated covariates for selection option $v$ be $\tilde{\mathbf{c}}_{vo}=(\tilde{c}_{ivo})_{i\in\mathcal{I}}$ and $\mathbf{s}_{vo}=(\mathbf{s}_{ivo})_{i\in\mathcal{I}}$, respectively. This paper assumes that option costs are independent, i.e.,
$\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})=\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo}),$
where $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ is the conditional distribution of the uncertain cost $\tilde{\mathbf{c}}_{vo}$ given the covariates $\mathbf{s}_{vo}$. This assumption is reasonable as the uncertain cost for a specific option depends solely on the contextual features of that option, not on the contextual features of other options. Note that, although these conditional distributions are considered independent, selection options that share common covariates (e.g., weather) and thus their distributions will tend to be similar. Problem (2) can then be rewritten as
$\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{c}}_{o}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})}[g(\mathbf{z}_{o},\mathbf{\tilde{c}}_{o})].$ (3)
3.3 Learning the Contextual Distribution
The conditional distribution $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is also generally unknown but, in practice, it can often be learned from historical data available. In many real-world settings however, each data point is observational, i.e., only those costs involved in the actual decision can be observed. As a result, the dataset available to estimate $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ is of the form $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{C}_{o})\}_{o=1}^{|\mathbb{D}|}$, where $\mathbf{C}_{o}$ are the observable costs for order $o$. The learning task consists in finding a machine learning $\mathcal{M}$ that approximates $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ in a supervised manner using the dataset $\mathbb{D}$: $\mathcal{M}$ receives a context $\mathbf{s}$ as input and returns a probability distribution $\mathcal{M}(\mathbf{s})$ of costs. Since learning $\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})$ amounts to learning $\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{vo})$ for each option $v$, the dataset $\mathbb{D}$ is partitioned into subsets, with $\mathbb{D}_{v}=\{(\mathbf{s}_{v1},\mathbf{c}_{v1}),(\mathbf{s}_{v2},\mathbf{c}_{v2}),\ldots\}$ representing the dataset for option $v$. A machine learning model $\mathcal{M}_{v}$ is associated with each option $v$ and $\mathcal{M}(\mathbf{\mathbf{s}})=(\mathcal{M}_{1}(\mathbf{s_{1}}),\ldots,\mathcal{M}_{|\mathcal{V}|}(\mathbf{s}_{|\mathcal{V}|}))$.
3.4 Contextual Sample Average Approximation (C-SAA)
Problem (3) is approximated by generalizing the Sample Average Approximation (SAA) method Kleywegt et al. (2002) to contextual stochastic optimization. The core of the Contextual Sample Average Approximation (C-SAA) for a request $o$ with covariates $\mathbf{s}_{o}$ amounts to
  1. 1.
    sampling the distribution $\mathcal{M}(\mathbf{s}_{o})$ $N$ times to obtain a set $\bm{\xi}_{o}^{n}=(\bm{\xi}_{vo}^{n})_{v\in\mathcal{V}}$ of cost vectors $(n\in[N])$;
  2. 2.
    solving the C-SAA problem
    (C-SAA) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n}).$ (4)
Following Kleywegt et al. (2002), the C-SAA procedure can be outlined as follows:
  1. 1.
    Solve $Q$ independent C-SAA problems, each of sample size $N=N_{1}$, to produce candidate solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ with corresponding objective values $\underaccent{\bar}{f}_{1},\ldots\underaccent{\bar}{f}_{Q}$. Calculate the average objective value as $\underaccent{\bar}{f}=\frac{1}{Q}\sum_{i=1}^{Q}\underaccent{\bar}{f}_{i}$. Let $f^{*}$ be the true optimal objective value of Problem (2). Since $\underaccent{\bar}{f}$ is an unbiased estimator of $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]$, and because $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s}_{o})}}[\underaccent{\bar}{f}]\leq f^{*}$, $\underaccent{\bar}{f}$ serves as a statistical lower bound on $f^{*}$ (Kleywegt et al. 2002).
  2. 2.
    Then, re-evaluate the solutions $\mathbf{z}_{1},\ldots,\mathbf{z}_{Q}$ using a larger sample size $N=N_{2}\gg N_{1}$, yielding new objective values $\bar{f}(\mathbf{z}_{1}),\ldots,\bar{f}(\mathbf{z}_{Q})$. Since each candidate solution is feasible, its re-estimated objective values $\bar{f}(\cdot)$ provides an upper bound on $f^{*}$. Therefore, the approximated upper bound on the optimality gap for each candidate solution is $\bar{f}(\cdot)-\underaccent{\bar}{f}$. Finally, select the best solution $\mathbf{z}^{*}$ from the candidates, which is the one with the smallest optimality gap.
It is interesting to compare C-SAA with the approach proposed in Bertsimas and Kallus (2020). A key idea in Bertsimas and Kallus (2020) is to use the historial data (the dataset $\mathbb{D}$ in this paper, as the samples for the stochastic optimization and to weight them properly. Unfortunately, this is not possible in this problem $\mathcal{P}$ considered here, because only observational data is available, i.e., only the costs of the selected options are revealed. One way to proceed would be to define the sample set for the cost vector as all possible combinations of data points across the selection options. However, as the number of selection options and historical data points increase, this approach suffers from the curse of dimensionality and becomes intractable for large dataset with big selection space, which is the case in the application considered in this paper.
3.5 Contextual Robust Optimization (C-RO)
The performance of C-SAA largely hinges on the accuracy of the estimated conditional distribution. To mitigate the risk of model misspecification in forecasting, robust optimization can be employed to account for worst-case scenarios, making it a viable approach for contextual stochastic optimization. Let $\mathcal{U}(\mathbf{s}_{o})$ represent the contextual uncertainty set for the cost vector $\tilde{\mathbf{c}}_{o}$. The Contextual Robust Optimization (C-RO) problem for (3) is then formulated as follows:
(C-RO) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\max_{\bm{\xi}_{o}\in\mathcal{U}(\mathbf{s}_{o})}g(\mathbf{z}_{o},\bm{\xi}_{o}).$ (5)
To effectively leverage the conditional distributional forecasts and create tractable reformulations of Problem (5), two types of uncertainty sets are considered: a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$ and a budgeted interval uncertainty set $\mathcal{U}_{b}(\mathbf{s}_{o})$.
3.5.1 Contextual Discrete Uncertainty Set
With a discrete uncertainty set $\mathcal{U}_{d}(\mathbf{s}_{o})$, Problem (5) can be reformulated into a single-level problem by enumerating all possible values of $\tilde{\mathbf{c}}_{o}$ within $\mathcal{U}_{d}(\mathbf{s}_{o})$. This is a common technique in robust optimization models (Goerigk and Khosravi 2024):
[…]
s.t. $t\geq g(\mathbf{z}_{o},\bm{\xi}_{o}),\quad\forall\bm{\xi}_{o}\in\mathcal{U}_{d}(\mathbf{s}_{o}).$
A discrete distribution can be generated as mentioned before, by combining all the datapoints for the selection options. However, this will make the problem intractable as mentioned earlier. For comparison purposes, the results will provide a sampling counterpart to that approach in the same spirit as C-SAA. Since it is not quite satisfactory, another uncertainty set is considered next.
3.5.2 Contextual Budgeted Interval Uncertainty Set
A commonly used uncertainty set is the budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$, which is defined by an uncertainty budget $B$. This budget limits the extent to which all components of the uncertain parameter $\tilde{\mathbf{c}}_{o}$ can simultaneously deviate to their upper bounds. This section assumes that the uncertain cost $\tilde{\mathbf{c}}_{o}$ is element-wise positive and $g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})$ is a piecewise linear function of the following form:
$g(\mathbf{z}_{o},\tilde{\mathbf{c}}_{o})=(\tilde{\mathbf{c}}_{o}+\mathbf{a}_{o}^{j})^{T}\mathbf{z}_{o}\quad\text{if}~{}\mathbf{z}_{o}\in[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}],~{}j=1,\ldots n,$ (7)
where $\{\mathbf{b}^{0}_{o},\mathbf{b}^{1}_{o},\ldots,\mathbf{b}^{n}_{o}\}$ are the breakpoints of the piecewise linear function, and $\mathbf{a}_{o}^{j}=(a_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}$ is a deterministic cost vector corresponding to the segment $[\mathbf{b}^{j-1}_{o},\mathbf{b}^{j}_{o}]=[(b_{ivo}^{j-1})_{i\in\mathcal{I},v\in\mathcal{V}},(b_{ivo}^{j})_{i\in\mathcal{I},v\in\mathcal{V}}]$. Given these assumptions, the typical approach to ensure tractability when using a budgeted interval set is to dualize the inner maximization problem, resulting in the robust counterpart of the original problem (Bertsimas and Sim 2003, Bayram et al. 2022, Goerigk and Khosravi 2024).
Let the conditional prediction interval of $\tilde{\mathbf{c}}_{vo}$ be $[\bm{\underline{\xi}}_{vo},\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}]$, where $\bm{\underline{\xi}}_{vo}=(\underline{\xi}_{ivo})_{i\in\mathcal{I}}$ and $\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo}=(\underline{\xi}_{ivo}+\hat{\xi}_{ivo})_{i\in\mathcal{I}}$ represent the lower and upper estimated conditional quantiles of the cost vector, respectively. The budgeted interval set $\mathcal{U}_{b}(\mathbf{s}_{o})$ can then be expressed as follows:
$\mathcal{U}_{b}(\mathbf{s}_{o},B)=\{\bm{\xi}_{o}\in\mathbb{R}^{|\mathcal{I}||\mathcal{V}|}:\xi_{ivo}=\underline{\xi}_{ivo}+\hat{\xi}_{ivo}\delta_{iv},\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B,\delta_{iv}\in[0,1]\}.$ (8)
For a budgeted uncertainty set, it should be noted that learning the entire conditional distribution is unnecessary. Instead, learning the lower and upper conditional quantiles is sufficient to construct the conditional prediction interval. For instance, to create a 95% conditional prediction interval, only the 2.5% and 97.5% conditional quantiles are required. Let $q^{low}$ and $q^{upp}$ be the lower and upper quantiles for the conditional prediction interval of selection option $v$, respectively. Then, the lower bound $\bm{\underline{\xi}}_{vo}$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{low}\}$, and the upper bound $(\bm{\underline{\xi}}_{vo}+\hat{\bm{\xi}}_{vo})$ is defined as $\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid{\mathbf{s}_{vo}})\geq q^{upp}\}$.
With a fixed $\mathbf{z}$, the inner maximization problem of Problem (5) is expressed as follows:
$\max\quad$ $\sum_{i\in\mathcal{I}}\sum_{v\in\mathcal{V}}\hat{\xi}_{ivo}z_{ivo}\delta_{iv}$ (9)
s.t. $\sum_{i\in\mathcal{I},v\in\mathcal{V}}\delta_{iv}\leq B$ (10)
$\delta_{iv}\in[0,1],\quad\forall i\in\mathcal{I},v\in\mathcal{V}.$ (11)
By taking the dual of the inner maximization problem and introducing auxiliary variables, a tractable reformulation of the robust counterpart of Problem (5) is obtained as follows:
[…]
s.t. $\pi+\lambda_{iv}\geq\hat{\xi}_{ivo}z_{ivo},\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
$\pi\geq 0$
$\lambda_{iv}\geq 0,\quad\forall i\in\mathcal{I},v\in\mathcal{V}$
$z_{ivo}\leq b^{j}_{ivo}+M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$z_{ivo}\geq b^{j-1}_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$w_{ivo}^{j}\geq(\underline{\xi}_{ivo}+a_{ivo}^{j})z_{ivo}-M(1-y_{j}),\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
$y_{j}\in\{0,1\},\quad\forall j\in[n]$
$w_{ivo}^{j}\geq 0,\quad\forall j\in[n],i\in\mathcal{I},v\in\mathcal{V}$
where $\pi$ and $(\lambda_{iv})_{i\in\mathcal{I},v\in\mathcal{V}}$ are the dual variables corresponding to Constraints (10) and (11), respectively, and $M$ is a reasonably large constant.
4 Application: The Ominichannel Multi-Courier Order Fulfillment Problem
This sections introduces a real-world order fulfillment problem encountered by the industrial partner. The proposed CSO framework is then applied to this problem, demonstrating significant advantages over the current practice and common baseline methods.
4.1 High-Level Problem Description
This section provides a high-level overview of the online order fulfillment challenges faced by the industrial partner. Similar to other major e-commerce companies, tens of thousands of orders are placed continuously throughout the day on the website and await processing. Each incoming order is organized by stock-keeping units (SKUs), resulting in one or more order lines. For each order line, there are hundreds of fulfillment centers across the company network eligible to source the SKUs. These centers include primarily physical stores, as well as regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping. The set of available carriers spans from major shipping companies to crowdsourced gig carriers. Figure 1 illustrates the order fulfillment life cycle within the operations of the industrial partner.
Figure 1: The Life Cycle of the Order Fulfillment Process.
Before making an order fulfillment decision, the company has access to both the transactional details of the order and the real-time status of the fulfillment network. Specifically, the order information includes the date and time the order was placed, the type, dimensions, weight, and quantities of the SKUs, the shipping destination, and the desired delivery date. The network status data provides details on the on-hand inventory levels of the requested SKUs and the remaining open capacities at various fulfillment centers. In addition, each carrier provides an estimated table for shipping costs and planned time-in-transit.
Given this information, the problem is to decide which fulfillment center to source each order line from and which carrier to use for shipping from that location. The goal is to identify the most cost-effective fulfillment option (a location-carrier pair) while ensuring the customer’s desired delivery date is met. The uncertainty of this problem lies in the delivery timeliness performance of each fulfillment option.
4.1.1 Overview of the Current Practice
Currently, the industrial partner uses a sourcing engine that employs a straightforward greedy algorithm. For each unit in an order, the sourcing engine sorts the eligible location-carrier pairs by the estimated shipping costs and transit times.
There are several limitations with this approach. First, since the engine makes decisions for each unit independently, it misses the opportunity to consolidate multiple units of the same or different products within an order. This can lead to higher shipping costs, as the potential to reduce the number of packages—and thereby lower shipping expenses—is not fully realized.
Second, the method relies solely on static transit time estimates provided by the carriers based on their service levels to select the best carrier for meeting the promised delivery date. This approach assumes that these estimates are always accurate. However, static transit times may not accurately reflect the actual delivery performance of carriers. Historical data from the industrial partner reveals that over 10% of orders experience varying degrees of early or late deliveries, which are called deviations in this paper. Given the complexity of the entire fulfillment process, these deviations can be attributed to numerous factors, such as warehouse pick-and-pack performance, late carrier arrivals, no-shows, and external factors like extreme weather or other unexpected circumstances. Such deviations from customer expectations can negatively impact their experience, potentially harming the company in the long run.
4.2 The Nominal Optimization Problem
This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem, which is a special case of the generic problem described in Section 3.1. A comprehensive list of the notations presented in this section is provided in Appendix for reference.
4.2.1 Main Decision Variables
Let $\mathcal{I}$ be the set of SKUs and $\mathcal{V}$ be the set of fulfillment options. Here, $\mathcal{V}$ is defined as the Cartesian product $\mathcal{K}\times\mathcal{L}$, where $\mathcal{K}=\{1,\ldots,K\}$ denotes the set of carriers, and =$\mathcal{L}=\{1,\ldots,L\}$ denotes the set of locations.
Each order $o$ needs to solve an independent optimization problem to decide the selections of carrier-location pairs for each SKU. The primary variables are denoted by the vector $\mathbf{z_{o}}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$, where each $z_{ik\ell o}\in\mathbb{Z}$ indicates the quantity of SKU $i$ sourced from location $\ell$ and shipped by carrier $k$ in order $o$.
4.2.2 Objective Function
The objective of the optimization problem for order $o$ is to minimize the overall fulfillment costs while ensuring timely delivery. The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair. Delivery timeliness is measured by deviations from the desired delivery date, whether early or late. To encourage timeliness, the objective imposes penalties for both early and late deliveries.
The objective function is formally defined as follows. Let $c_{ik\ell o}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$ in order $o$. For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost. For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center. Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
In an ideal setting where all carrier-location deviations are known at the time of fulfillment, let the vector $\mathbf{d}_{o}=(d_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ represent the realized uncertain delivery deviations for order $o$, where each $d_{k\ell o}\in\mathbb{R}$ denotes the number of days that the carrier-location pair $(k,\ell)$ deviates from desired delivery date. A positive deviation indicates late delivery, a zero deviation indicates on-time delivery, and a negative deviation indicates early delivery.
The weighted objective function for order $o$ can be constructed as follows:
$g(\mathbf{z}_{o},\mathbf{d}_{o})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-}]z_{ik\ell o}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell o}\geq 2\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship}+\gamma^{+}d_{k\ell o}^{+}+\gamma^{-}d_{k\ell o}^{-})z_{ik\ell o}&\text{otherwise},\end{cases}$ (13)
where $d_{k\ell o}^{+}=\max\{0,d_{k\ell o}\}$ and $d_{k\ell o}^{-}=\max\{0,-d_{k\ell o}\}$. This objective function encourages item consolidation within an order by offering a carrier-specific discount when applicable. In addition, it leverages an asymmetric penalty cost function for delivery deviations, with $\gamma^{+}\gg\gamma^{-}$ to primarily discourage late deliveries while also disincentivizing early deliveries to a lesser extent. The objection function is piecewise linear and can be linearized using standard rewritings.
4.2.3 Operational Constraints
Several operational constraints ensure that the selection of fulfillment center locations and carriers is feasible in practice. For location $\ell\in\mathcal{L}$, let $\operatorname{inv}_{i\ell o}\geq 0$ be the inventory level of SKU $i$ at location $\ell$ and $\operatorname{cap}_{\ell o}\geq 0$ be the available capacity at location $\ell$ when order $o$ is placed. For SKU $i$ in order $o$, let $q_{io}\geq 0$ be the quantity of SKU $i$ in order $o$, and let $e_{ik\ell o}\in\{0,1\}$ be a binary indicator of whether the location-carrier pair $(\ell,k)$ is an eligible fulfillment option.
  • Each unit of an SKU must be sourced from exactly one location and shipped by exactly one carrier.
    $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}z_{ik\ell o}=q_{io},\quad\forall i\in\mathcal{I}$ (14)
  • Each SKU can only be fulfilled by eligible candidate location-carrier pairs, which may vary depending on the carrier’s coverage and package requirements.
    $z_{ik\ell o}\leq q_{io}e_{ik\ell o},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$ (15)
  • Only locations with available inventory of the requested SKUs can be chosen.
    $\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{inv}_{i\ell o},\quad\forall i\in\mathcal{I},\ell\in\mathcal{L}$ (16)
  • Each location has a daily capacity limit on the total number of units it can process.
    $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}z_{ik\ell o}\leq\operatorname{cap}_{\ell o},\quad\forall\ell\in\mathcal{L}$ (17)
4.3 Applying the CSO Framework to the Problem
In practice, the delivery deviations of each carrier-location pair cannot be perfectly estimated in advance. To account for this uncertainty for each order $o$, this section models the vector of uncertain delivery deviations $\tilde{\mathbf{d}}_{o}=(\tilde{d}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}\sim\mathbb{P}$ as a $(KL)$-dimensional random variable governed by the probability distribution $\mathbb{P}$. In addition, let $\mathbf{s}_{o}=(\mathbf{s}_{k\ell o})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ denote the state vector observable for order $o$ before the fulfillment decision is made, where each $\mathbf{s}_{k\ell o}$ is a vector of covariates encompassing information about the fulfillment network status and order details related to carrier-location pair $(k,\ell)$. The conditional distribution of the deviation vector, upon observing $\mathbf{s}_{o}$, is denoted by $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$, and the individual deviation condition distribution is $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ for each $k\in\mathcal{K}$ and $\ell\in\mathcal{L}$.
Applying the CSO framework leads to the following contextual stochastic order fulfillment problem (CSOFP):
(CSOFP) $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\Pi_{k\in\mathcal{K},\ell\in\mathcal{L}}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})].$ (18)
Effectively solving Problem (18) requires leveraging data from historical orders. As a specific instance of Problem (2), CSOFP also has an observational dataset. Let $\mathcal{O}$ be the set of historical orders. Let $\mathbf{z}_{o}=(z_{ik\ell o})_{i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}}$ be the vector of historical decisions made by the industrial partner’s sourcing engine for order $o$. There are numerous missing entries in the observed deviation vector $\mathbf{d}_{o}$ since actual observations depend on the decision vector $\mathbf{z}_{o}$. Let
$\mathbf{D}_{o}:=\{d_{k\ell o}\mid\sum_{i\in\mathcal{I}}z_{ik\ell o}>0,k\in\mathcal{K},\ell\in\mathcal{L}\}$
be the observable costs for order $o$. The dataset available to estimate $\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})$ is $\mathbb{D}=\{(\mathbf{s}_{o},\mathbf{D}_{o})\}_{o=1}^{|\mathcal{O}|}$.
$\mathbb{D}$ is partitioned into subsets to facilitate learning the individual condition distributions. For carrier $k$ and location $\ell$, the dataset for learning $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell o})$ is $\mathbb{D}_{k\ell}=\{(\mathbf{s}_{k\ell m},d_{k\ell m})\}_{m=1}^{M_{k\ell}}$, where $M_{k\ell}$ is the number of times the deviation has been observed for the carrier-location pair $(k,\ell)$ in $\mathbb{D}$.
4.4 Learning the Contextual Distribution of Delivery Deviations
This section describes how to learn Contextual Distribution of Delivery Deviations. It begins by selecting features to construct the covariate vector for each carrier-location pair. To predict the conditional distribution of deviations, two classes of machine learning methods are proposed: probabilistic multi-class classification, which leverages the discrete nature of deviations, and quantile regression, which is known for its effectiveness in distributional prediction (Salari et al. 2022). For simplicity, the index $k\ell$ is omitted in the notations used throughout this section. The data set can then be denoted by $\mathbb{D}=\{(\mathbf{s}_{m},d_{m})\}_{m\in[M]}$.
4.4.1 Contextual Features Selection
Carefully selecting contextual features is crucial for constructing informed covariates for delivery deviation predictions, as various stages in the order life cycle can contribute to deviations. To identify the most important features for predicting deviations, a feature importance analysis was conducted using a tree-based mean regression model. Two methods were examined: Mean Decrease in Impurity (MDI) and Shapley values (SHAP). MDI measures the importance of a feature by calculating the total reduction in impurity (e.g., Gini impurity or entropy) it brings when used in decision trees (Breiman 2001). SHAP values, on the other hand, provide a unified measure of feature importance based on cooperative game theory, attributing to each feature its contribution to the prediction by considering all possible feature combinations (Lundberg et al. 2020).
The features considered are categorized into three main classes, each corresponding to different stages of the order fulfillment process:
  1. 1.
    Order-level features: This category includes customer location, SKU weight and dimensions, SKU type, quantities, release hour, release day, and planned lead time (i.e., the difference between the promised delivery date and the order release time). These factors can influence deviations through logistical complexity, handling requirements, and time-sensitive elements affecting the order’s processing and delivery schedule.
  2. 2.
    Fulfillment center features: This includes coordinates, location type, on-hand capacity, and on-hand inventory. The geographical location and type of fulfillment center can affect the efficiency of order processing, while capacity and inventory levels determine the speed and reliability of fulfilling orders.
  3. 3.
    Carrier-related features: This includes the level of service, promised time in transit, shipping charges, and carrier zone. These factors impact the carrier ability to meet delivery deadlines, with service level and transit times directly affecting deviations, while shipping charges and zones reflect logistical challenges and potential bottlenecks.
4.4.2 Probabilistic Multi-Class Classification
Since deviations measured in days only take on discrete values in the dataset, they inherently follow a discrete probability distribution. Thus, the random variable $\tilde{d}$ has a finite discrete support. Let $C$ be the total number of possible values $\tilde{d}$ can take, and let $\xi_{c}$ be the $c$-th possible value for $c\in[C]$. By definition, $\tilde{d}\in\{\xi_{1},\xi_{2},\ldots,\xi_{C}\}$. This discrete nature allows the conditional distribution learning problem to be framed as a multi-class classification (MC-CLF) problem. Under MC-CLF, each class represents a possible delivery deviation value in number of days, and the output class probabilities correspond directly to the probability mass function of the discrete distribution. For a new sample with covariates vector $\mathbf{s}$, let the conditional probability of class $c$ be $p_{c}(\mathbf{s})=\mathbb{P}(d=\xi_{c}\mid\mathbf{s})$, with the corresponding predicted probability from MC-CLF denoted as $\hat{p}_{c}(\mathbf{s})$.
Multinomial logistic regression (MLR) and classification trees-based models are considered to solve MC-CLF. Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, which is the negative of the log loss function. The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree. Given covariates $\mathbf{s}$, the predicted probability for class $c$ in a classification tree can be expressed as:
$\hat{p}_{c}(\mathbf{s})=\frac{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}\}\cdot\mathbf{1}\{d=\xi_{c}\}}{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}\}},$ (19)
To enhance predictive power, ensemble learning methods such as bagging and gradient boosting are employed, through the use of random forests (RF) and gradient boosted trees (GBT). The details for their calculation of the class probabilities are available in Appendix.
Moreover, while the proposed classification models excel at predicting class labels, their predicted probabilities may not be well calibrated (Niculescu-Mizil and Caruana 2005). Therefore, probability calibration is necessary to obtain accurate conditional distributional forecasts. To address this, a cross-validation approach is adopted to fit an isotonic regression calibrator (Zadrozny and Elkan 2002). In each split, the classification model is first trained on the training set, and then its predictions on the calibration set are used to fit an isotonic regression model. Let $\hat{\tilde{p}}(\cdot)$ be the vector of calibrated probabilities and $\hat{p}(\cdot)$ be the vector of uncalibrated probabilities for each class. Isotonic regression takes the vector of uncalibrated probabilities as input and solves the following problem to determine a vector of monotonically increasing calibrated probabilities:
$\min_{\hat{\tilde{p}}(\cdot)}$ $\sum_{m=1}^{M}(d_{m}-\hat{\tilde{p}}(\mathbf{s}_{m}))^{2}$ (20)
s.t. $\hat{\tilde{p}}(\mathbf{s}_{m})\geq\hat{\tilde{p}}(\mathbf{s}_{m^{\prime}})\quad\forall m^{\prime}\neq m:\hat{p}(\mathbf{s}_{m})\geq\hat{p}(\mathbf{s}_{m^{\prime}})$
To measure the accuracy of probabilistic predictions for MC-CLF model, the paper adopts the log loss and the Brier score. The latter is defined as the average squared error between predicted probabilities and actual outcomes $\sum_{m=1}^{M}\sum_{c=1}^{C}(\mathbf{1}\{d_{m}=\xi_{c}\}\hat{p}_{c}(\mathbf{s}_{m}))^{2}$.
For both metrics, lower values indicate better probabilistic prediction performance.
4.4.3 Quantile Regression
Unlike multi-class classification, quantile regression (QR) can directly estimate the cumulative distribution function (CDF) of the conditional distribution without assuming a discrete probability distribution. The CDF of the uncertain deviation $\tilde{d}$ is given by $F(u\mid\mathbf{s})=\mathbb{P}(d\leq u\mid\mathbf{s})$. Standard QR models aim at minimizing the pinball loss function, estimating different quantiles independently. To approximate the conditional distribution, these models typically calculate a series of conditional quantiles, forming a discretized approximation of the CDF. However, this approach can lead to the quantile crossing issue, where predicted quantiles do not maintain their proper order, resulting in an invalid distribution. Various methods exist to avoid quantile crossing, but tree-based quantile regression models have been shown to outperform standard QR models in distributional forecasting, particularly in delivery time performance settings (Salari et al. 2022). These tree-based methods do not suffer from the issue of quantile crossing. Therefore, this paper focuses on a regression tree-based model and a quantile random forest (QRF) model which leverages an ensemble of regression trees.
Similar to a classification tree, a regression tree partitions the feature space into multiple disjoint regions, but the split criterion is based on the mean squared error (MSE). While regression trees are typically used for generating conditional point predictions, they can be extended to provide distributional forecasts by preserving all observations within each partitioned region, rather than just summary statistics like the mean. Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a regression tree. Given covariates $\mathbf{s}$, the region that $\mathbf{s}$ falls into is denoted as $\mathcal{R}(\mathbf{s})$. The conditional distribution of the new sample can be approximated by the empirical CDF:
$\hat{F}(u\mid\mathbf{s})=\sum_{m=1}^{M}\frac{\mathbf{1}\{\mathbf{s}_{m}\in\mathcal{R}(\mathbf{s})\}}{|\{j:\mathbf{s}_{j}\in\mathcal{R}(\mathbf{s})\}|}\mathbf{1}\{d_{m}\leq u\}.$ (21)
To refine this empirical CDF, a kernel density estimation method is applied to the observations in $\mathcal{R}(\mathbf{s})$ (Salari et al. 2022).
A similar approach can be extended to random forests, and this idea of QRF was first introduced by Meinshausen and Ridgeway (2006). In QRF, for a new sample with covariates $\mathbf{s}$, the empirical distribution is derived from all regions associated with $\mathbf{s}$ across all trees in the forest. Let $T$ be the total number of regression trees, and let $\mathcal{R}_{1}^{(t)},\ldots,\mathcal{R}_{R}^{(t)}$ be the partitioned regions of the $t$-th tree, where $t\in[T]$. The region that $\mathbf{s}$ falls into in the $t$-th tree is denoted as $\mathcal{R}^{(t)}(\mathbf{s})$. QRF approximates the conditional distribution using the empirical CDF:
$\hat{F}(u\mid\mathbf{s})=\frac{1}{T}\sum_{t=1}^{T}\sum_{m=1}^{M}\frac{\mathbf{1}\{\mathbf{s}_{m}\in\mathcal{R}^{(t)}(\mathbf{s})\}}{|\{j:\mathbf{s}_{j}\in\mathcal{R}^{(t)}(\mathbf{s})\}|}\mathbf{1}\{d_{m}\leq u\}.$ (22)
To measure the distributional forecasting performance of different learning models, the paper uses two metrics—the Continuous Ranked Probability Score (CRPS) and pinball loss. CRPS evaluates the closeness of the estimated CDF $\hat{F}$ given covariates $\mathbf{s}$ to the observed ground truth value $d$:
[…]
Let $\xi$ and $\xi^{\prime}$ be two independent random variables with distribution $\hat{F}(u\mid\mathbf{s})$. The CRPS can alternatively be expressed as:
$\text{CRPS}(\hat{F},d,\mathbf{s})=\mathbb{E}(|\xi-d|)-\frac{1}{2}\mathbb{E}(|\xi-\xi^{\prime}|).$ (24)
This equation allows for an approximation of CRPS by sampling from $\hat{F}(u\mid\mathbf{s})$. Pinball loss measures the accuracy of quantile predictions by penalizing errors asymmetrically, depending on whether the predictions are overestimates or underestimates. Let $q$ be a specific quantile. Then, the pinball loss between the observed value $d$ and the predicted value $\hat{d}$ is defined as:
$\text{pinball}(d,\hat{d})=q\max\{d-\hat{d},0\}+(1-q)\max\{\hat{d}-d,0\}.$ (25)
4.4.4 Applying the C-SAA and the C-RO
It is impractical to solve the CSOFP exactly with the learning models just proposed. Indeed, on the one hand, QR models learn continuous distributions, making it computationally intractable to expand the expectation directly. On the other hand, MC-CLF models leverage the fact that the uncertain delivery deviation vector $\tilde{\mathbf{d}}$ has a finite discrete support, which theoretically allows for expansion of the expectation term. However, MC-CLF models face the curse of dimensionality. The number of possible values of $\tilde{\mathbf{d}}$ grow exponentially in $K$ and $L$. This justifies the methodologies proposed in Section 3.
In the following, the C-SAA problems using samples from the conditional distribution estimates provided by the MC-CLF and QR models are referred to as C-SAA-CLF and C-SAA-QR, respectively. Similarly, the C-RO problems are referred to as C-RO-{B,D}-CLF and C-RO-{B,D}-QR, respectively.
5 Case Study
This section presents computational results that demonstrate the potential of the proposed CSO framework when evaluated on real-world industrial instances.
5.1 Data Overview
The industrial partner provides a month-long dataset in 2023, covering more than 1 million online orders for home delivery and over 20,000 unique stock-keeping units (SKUs). Each order is organized by SKUs, resulting in one or more order lines. For an incoming order, there are, on average, 250 to 1,000 possible fulfillment centers across the company’s network to source SKUs from, including mostly physical stores but also regional distribution centers and metro e-commerce centers. Each location is equipped with multiple eligible carriers that are capable of providing different levels of service (LOS), such as Ground, Same-Day, and Overnight shipping. The set of available carriers span from major shipping companies to crowdsourced gig carriers. There are 10 to 30 carrier-LOS pairs appeared in the dataset and throughout the paper, “carrier” refers to one of these specific carrier-LOS pairs.
Each row of the dataset contains transactional details for a single order line, along with the status of the fulfillment network at the time the order was placed. The order information includes the date and time of order placement, the type, dimensions, weight, and quantity of the SKUs, the fulfillment location from which each order line was sourced, the carrier that shipped the order, the customer location, the promised delivery date, and the actual arrival date. The network status data includes the on-hand inventory levels of the requested SKU and the remaining open capacity at various locations. Additionally, the dataset includes carrier-provided shipping costs and planned transit times.
5.1.1 Learning Results
For the evaluation of the contextual learning models, the dataset is split into two parts: orders from the first three weeks (1.3 million order lines) are used for training, while orders from the last week (0.4 million order lines) are reserved for testing and evaluation.
In Table 1, two metrics—log loss and Brier score—are reported to measure the accuracy of probabilistic predictions for the tuned classification models on the test set. Probability calibration is applied to the uncalibrated models that obtained the best metrics. It is notable that both the log loss and Brier scores show improvement after calibration. In addition, Figure 2 illustrates the performance of an example distributional forecast by multi-class classification against its true empirical distribution on the test set. The distribution of the majority class 0 (on-time deliveries) is predicted quite accurately, while the distributions of late and early deliveries are forecasted sufficiently close, albeit with small deviations.
Table 1: Average Log Loss and Brier Score on Test Set
Model MLR Classification Tree RF-CLF GBT-CLF Calibrated-CLF
Log loss 1.950 2.242 1.593 1.050 0.965
Brier score 0.073 0.072 0.054 0.034 0.031
Figure 2: Comparison of Calibrated Predicted Joint Distribution with True Empirical Distribution for an Example Carrier on Test Set.
Table 2 presents the pinball losses for different quantiles and the CRPS values of the tuned quantile regression models and classification models. The CRPS values are approximated by drawing 1,000 samples to compute Equation (24). It is observed that the MC-CLF models exhibit higher pinball losses compared to the two quantile regression models, likely due to the discrete nature of classification outputs. However, the CRPS of MC-CLF is the lowest among the three models, despite not being explicitly tuned for CRPS.
Additionally, Figure 3 displays the 95% prediction intervals created by the three conditional distribution learning methods for an example carrier. The visualization aligns with the pinball loss results, showing that QRF generally provides the narrowest interval widths. While the intervals generated by MC-CLF cover most of the observed values, their coverage is less effective at the two extremes compared to the two quantile regression methods. The inferior performance of MC-CLF at extreme quantiles is likely due to the discretization of the output space, which limits precision, and the cumulative probability method, which can cause jumps when selecting quantiles. In contrast, regression tree and QRF provide finer-grained predictions, leading to more accurate and narrower intervals at the extremes.
In summary, MC-CLF models excel at providing distributional forecasts, whereas QR models are better at generating prediction intervals for this dataset.
Table 2: Average Pinball Loss and CRPS on Test Set Across All Carriers.
Pinball loss CRPS
Model $q_{0.025}$ $q_{0.05}$ $q_{0.5}$ $q_{0.95}$ $q_{0.975}$
MC-CLF 0.199 0.198 0.676 0.526 0.304 0.386
Regression Tree 0.064 0.109 0.302 0.147 0.093 0.515
QRF 0.033 0.057 0.263 0.143 0.096 0.402
(a) Regression Tree Prediction Intervals
(b) MC-CLF Prediction Intervals
(c) QRF Prediction Intervals
Figure 3: Sorted Observed Deviation Values versus 95% Prediction Intervals for an Example Carrier.
5.2 Baseline Models
To compare with the proposed contextual optimization models, three baseline order fulfillment algorithms are introduced below.
  1. 1.
    Pure Cost-Driven Greedy Heuristic (Greedy): This greedy heuristic replicates the current practice employed by the industrial partner’s sourcing engine. The algorithm selects the location-carrier pair with the lowest fulfillment cost for each quantity of an SKU. For order $o$, the greedy problem is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell o}^{ship})z_{ik\ell o}\}.$
  2. 2.
    Empirical-SAA: This method follows the approach described in Section 3.4 for C-SAA, with the key difference being that the deviation vector samples are drawn from the empirical distribution of deviations in the training set, without incorporating any contextual information. For each sample $n\in[N]$, let $\xi_{k\ell o}^{n}$ denote the $n$-th devaition value sampled from the empirical distribution $P^{E}_{k\ell}$ for carrier-location pair $(k,\ell)$. The corresponding deviation vector for sample $n$ is then expressed as $\bm{\xi}_{o}^{n}=(\xi_{k\ell o}^{n})_{k\in\mathcal{K},\ell\in\mathcal{L}}$. For order $o$, the empirical SAA problem is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}\{\frac{1}{N}\sum_{n=1}^{N}g(\mathbf{z}_{o},\bm{\xi}_{o}^{n})\}.$
  3. 3.
    Point Predict, Then Optimize (Point-Pred.-Opt.): Unlike the distributional forecasting models mentioned in Section 4.4, typical regression-based ML models aim at estimates only the conditional mean $\mathbb{E}[\tilde{d}_{k\ell}\mid\mathbf{s}_{k\ell o}]$. In this method, conditional mean regression models $h_{k\ell}$ are trained to map covariates to deviation for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$. This approach transforms CSOFP into a deterministic optimization problem by directly replacing the random parameter $\tilde{\mathbf{d}}_{o}$ with its point prediction $\hat{\mathbf{d}}_{o}=\{h_{k\ell}(\mathbf{s}_{k\ell o})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$ in the objective function of Problem (18). This method is effective only if the objective function $g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})$ is linear in $\tilde{\mathbf{d}}_{o}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean. This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}}_{o})]=g(\mathbf{z}_{o},\mathbb{E}_{\tilde{\mathbf{d}}_{o}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s}_{o})}[\tilde{\mathbf{d}}_{o}])$. However, this linearity does not hold in the proposed objective function due to the asymmetric penalty costs associated with the deviations. So a point prediction may be biased, potentially leading to suboptimal solutions. To facilitate the experiments, several different regression models are trained to minimize MSE, and the model with the smallest MSE in cross-validation is selected for each carrier. See Appendix for the average MSE of different regression models. For order $o$ and point prediction $\hat{\mathbf{d}}_{o}$, the optimization model is
    $\min_{\mathbf{z}_{o}\in\mathcal{Z}_{o}}g(\mathbf{z}_{o},\hat{\mathbf{d}}_{o}).$
5.3 The Experimental Setting
The optimization models in this study were implemented using Gurobi 11.0.0 and Python 3.9. All experiments were conducted on a server equipped with dual-socket Intel Xeon Gold 6226 CPUs, each featuring 24 cores running at 2.7 GHz. The regression and multi-class classification models were developed using the scikit-learn (Pedregosa et al. 2011), XGBoost (Chen and Guestrin 2016), and CatBoost libraries (Prokhorenkova et al. 2018). For quantile regression, the regression tree and quantile random forest models were implemented using the quantile-forest package (Johnson 2024).
For the following experiments, all ML models are trained on orders from the first three weeks, while 5,000 multi-item orders randomly sampled from the last week are used to evaluate the order fulfillment methods. During both training and inference, for multi-SKU orders, features are aggregated by retaining only the features of the SKU with the highest shipping charges and quantities.
One of the difficulties in this application is that the ground truth is not available for options that were not selected. To overcome this limitation, a simulation environment was built and generate the “ground truth” when an algorithm selects an option. The delivery deviations are drawn from a multinomial distribution whose probabilities were estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders in the last week of the dataset. Obviously, this data is not available to the algorithm and only used by the simulator to replicate real-world conditions.
By default, a constant $\beta_{k}=0.5$ discount factor for item batching is considered since the exact values are not provided. Sensitivity analysis are perform on the impact of this factor. The default penalty for early delivery $\gamma^{-}$ is $0.1$ and that for late delivery $\gamma^{+}$ is $10$. The penalty can be understood as the opportunity cost (lost sale cost) for early (late) deliveries, and they are also analyzed with a range of different values. For methods relying on SAA, the default is to have $Q=20$ independent replications each with a sample size of $N_{1}=30$. The obtained solutions are then evaluated with a default sample size of $N_{2}=300$. The default sample size used for C-RO-D is $N=200$. As for C-RO-B, a default of $95\%$ interval is considered. Moreover, a scaling factor $\eta\in[0,1]$ (default of 0.5) determines the uncertainty budget, defined as $B=\eta(KL).$
5.4 Effectiveness of C-SAA
This section first compares the performance of the C-SAA method against the three baseline methods. Throughout the following sections, only normalized values of the metrics are reported due to the confidentiality of the data.
5.4.1 Overall Results
Table 3 compares the realized objective value of Problem (18) for C-SAA and the baseline methods. For C-SAA, only the metrics of the best prediction model are recorded. It can be observed that C-SAA consistently has the lowest objective value across different late delivery penalties. C-SAA maintains a lower objective value on average across all $\gamma^{+}$ levels (approximately 0.80), compared to Point-Pred.-Opt. (0.87) and Empirical-SAA (0.83), highlighting its stability and reliability.
A perfect balance between the fulfillment costs and delivery timeliness control needs to be tailored for the specific business needs. On the one hand, a lower fulfillment cost brings mores immediate profits. On the other hand, lower delivery timeliness can significantly increase customer satisfaction, thus retaining more loyal customers, benefiting the company in long-term operations. In CSOFP, the trade-off between these short-term and long-term profits is controlled by the key parameter $\gamma^{+}$, which determines the penalty of late deliveries in terms of costs. The trend indicates that C-SAA’s objective value decreases more sharply compared to other methods as $\gamma^{+}$ increases, making it particularly effective in scenarios where delivery timeliness is crucial. C-SAA outperforms Point-Pred.-Opt. by 2% (0.92 vs. 0.94) when $\gamma^{+}=2.5$ and by 14% (0.63 vs. 0.79) when $\gamma^{+}=40$, indicating that C-SAA’s advantage becomes more pronounced as the penalty increases. While the difference between C-SAA and Empirical SAA in terms of objective value is not huge, the subsequent section will further illustrate the performance gaps between these two methods, as Empirical SAA is more sensitive and could potentially perform worse with larger sample sizes.
Table 3: Simulated Realized Objective Values of C-SAA vs. the Three Baselines
Method $\gamma^{+}=2.5$ $\gamma^{+}=5$ $\gamma^{+}=10$ $\gamma^{+}=20$ $\gamma^{+}=40$
Greedy 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-) 1.0 (-)
Point-Pred.-Opt. 0.94 $(2.6\times 10^{-4})$ 0.92 $(3.2\times 10^{-3})$ 0.88 $(4.2\times 10^{-3})$ 0.83 $(5.5\times 10^{-3})$ 0.79 $(7.0\times 10^{-3})$
Empirical-SAA 0.93 $(0.9\times 10^{-4})$ 0.91 $(2.0\times 10^{-3})$ 0.86 $(3.1\times 10^{-3})$ 0.76 $(1.0\times 10^{-2})$ 0.64 $(1.2\times 10^{-2})$
C-SAA 0.92 $(1.3\times 10^{-4})$ 0.89 $(3.3\times 10^{-3})$ 0.82 $(5.5\times 10^{-3})$ 0.72 $(1.1\times 10^{-2})$ 0.63 $(1.3\times 10^{-2})$
  • Note: Values represent the normalized mean (standard error) across five different random seeds for the realized deviations simulation. The normalization is performed by setting the value for the Greedy method to 1.
5.4.2 Value of Contextual Information
To facilitate a closer examination of the performance of these SAA-based models, Figure 4 shows the impact of varying sample sizes. On the one hand, larger sample sizes should provide a more accurate approximation of the expected value in Problem (18), assuming the delivery deviation distributions are well-estimated. On the other hand, it is important to maintain a reasonably small sample size to ensure computational efficiency. For both C-SAA methods, the effect of increasing sample sizes is obvious. The objective values are reduced from 0.86 to 0.82 for C-SAA-CLF and 0.87 to 0.84 for C-SAA-QR on average, indicating their increasing alignment with the true risk-neutral objective function. This improvement is particularly evident in their enhanced effectiveness at avoiding cumulative lateness, defined to be the total number of late delivery days across all units in the orders. The cumulative lateness decreases by 21% for C-SAA-CLF and by 18% for C-SAA-QR.
In contrast, Empirical-SAA struggles to improve its performance in both the realized objective value and the cumulative lateness as the sample sizes grows beyond 10. This can be attributed to the fact that purely using the empirical distribution is not sufficient to capture the varying delivery deviation distributions that may be dependent upon contextual information unique upon each order. A biased distributional forecasts thus guide the model wrongly to create solutions with inferior quality. Based on these observations, both conditional distributional forecasting methods proposed in the paper are successfully in this application, demonstrating a stronger performance compared to using empirical distribution for delivery deviation forecasting. In summary, the value of contextual information is particularly important in order fulfillment decision-making, where the uncertainty of delivery deviation can vary drastically across different states of the environment.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 4: C-SAA vs. Empirical SAA Across Different Sample Sizes.
5.4.3 Value of Distributional Forecasts
This section examines the advantage of employing distributional forecasting models over the simpler and more common point prediction model for CSOFP. As supported by theory, simply using the conditional mean estimate of deviations to replace the expected value in CSOFP can introduce bias when approximating the true objective value. This approach may yield worse solutions compared to employing conditional distributional estimators. These performance differences become even more pronounced as late penalties increase, as illustrated in Figure 5. With higher late penalties, accurate forecasting of deviations becomes more crucial; the results reveal the inherent bias of Point-Pred.-Opt, leading to lower quality solutions. This is not only evident in the realized objective value (Figure 5(a)) but also in the cumulative lateness (Figure 5(b)). For example, the cumulative lateness for Point-Pred.-Opt. only improves by 7% from $\gamma^{+}=2.5$ to $\gamma^{+}=40$, whereas C-SAA-CLF and C-SAA-QR show significantly greater improvements of approximately 35% and 32%, respectively. This demonstrates that the C-SAA methods not only maintain lower cumulative lateness overall but also improve more significantly as the penalty for late deliveries increases, further highlighting the benefits of distributional forecasting.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 5: C-SAA vs. Point-Pred.-Opt. Across Different Late Delivery Penalties.
5.5 Evaluation of C-RO
This section evaluates the performance of the C-RO methods, whose results are shown in Figure 6. The C-RO methods generally provide more conservative solutions than C-SAA, with C-RO-B showing a performance closer to C-SAA, as shown by the realized objective values in Figure 6(a). Notably, C-RO-B outperforms C-RO-D by offering less conservative yet still robust solutions, while not requiring full distributional information. In real-time fulfillment scenarios, where sampling can be time-prohibitive, especially during peak hours, C-RO-B offers a quick and effective alternative, delivering reasonably high-quality solutions. Additionally, the effectiveness of C-RO methods in controlling late deliveries remains consistent as the penalty increases (see Figure 6(b)). This stability occurs because C-RO methods already account for worst-case scenarios, making them less sensitive to changes in the late delivery penalty, provided that the samples or uncertainty sets encompass these scenarios. As a result, C-RO methods outperform C-SAA in reducing cumulative lateness when the late delivery penalty is low but are outperformed by C-SAA when $\gamma^{+}>10$. However, the C-RO methods incur higher fulfillment costs—calculated as the total shipping costs and fixed non-parcel costs, minus item consolidation discounts across all orders—by up to 20% on average (see Figure 6(c)) This increase is driven by their risk-averse approach, which prioritizes avoiding extreme deviations by opting for more expensive fulfillment options.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
(c) Performance in Fulfillment costs.
Figure 6: Performance of C-RO Methods.
5.5.1 Flexibility of C-RO-B
The effectiveness of the C-RO models in controlling delivery deviations is evident since they can hedge against the worst-case deviation and potential misspecification of the prediction models. However, their robustness to extreme delivery deviations comes at the cost of more conservative solutions, which can result in higher fulfillment expenses. Fortunately, the C-RO-B models, beyond their distribution-free property, are equipped with the flexibility to control their level of conservativeness. Specifically, they are determined via the coverage of the prediction intervals and the robust uncertainty budget. Smaller prediction interval coverage and uncertainty budget may lead to less conservative solutions, making the models more beneficial in terms of addressing forecasting misspecifications rather than purely managing risk. Conversely, larger coverage and budget aim at minimizing the impact of worst deviations as much as possible with the sacrifice of overall solution quality.
The impact of different prediction interval coverages and robust uncertainty budgets on the quality of solutions obtained from C-RO-B is examined in Figures 7 and 8, respectively. As the prediction interval coverage increases from 50% to 95% in Figure 7(a) , both C-RO-B-CLF and C-RO-B-QR show a noticeable decline in the realized objective value, with the lowest values around 0.85 for C-RO-B-CLF and 0.86 for C-RO-B-QR at 95% and 90% coverage, respectively. This represents a reduction of approximately 6.5% for both C-RO-B methods compared to the objective values at 50% coverage. This suggests that increasing the coverage interval initially benefits the overall performance, up to a point. This is consistent with the general decline trend in cumulative lateness as prediction interval increases (see Figure 7(b)). However, at 99% coverage, there is a sharp increase in the objective value, This trend highlights that while conservative prediction intervals can reduce lateness, they also introduce higher fulfillment costs due to more conservative decisions, thus increasing the overall objective value. For C-RO-B-CLF, this increase in objective value is also due to a declined prediction performance with the largest prediction interval, as indicated by the final increase of cumulative lateness at 99% coverage.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 7: Performance of C-RO-B Across Different Prediction Interval Coverages.
A similar pattern is observed for varying robust budgets. As $\eta$ increases, the robust budget becomes larger, leading to more conservative solutions. For both C-RO-B methods, larger budget results in a more significant decrease in cumulative lateness (see Figure 8(b)), but the results are less clear when it comes to the objective value (see Figure 8(a)). The realized objective value decreases sharply as the robust budget scaling factor $\eta$ increases from 0.0 to 0.01 for both C-RO-B-CLF and C-RO-B-QR. However, after $\eta=0.01$, there is a noticeable jump in the objective value for C-RO-B-QR, while C-RO-B-CLF continues to stabilize. This jump suggests that C-RO-B-QR adopts more conservative decisions with higher robust budgets, which leads to an increase in fulfillment costs. Therefore, C-RO-B-QR hits its lowest objective value at $\eta=0.01$ whereas C-RO-B-CLF is able to obtain its minimum objective value at $\eta=1$. The analysis reveals a trade-off between increasing robustness and striking a balance between lower fulfillment costs and mitigating extreme deviation risks.
(a) Performance in Realized Objective Value.
(b) Performance in Cumulative Lateness.
Figure 8: Performance of C-RO-B Across Different Robust Uncertainty Budgets.
5.6 Impact of Consolidation Factor
This section provides a discussion on the impact of item consolidation. CSOFP relies on the coefficients $\beta_{k}$ to control the incentive to batch items. The industrial partner does not know the exact percentage of discount to be offered for item consolidation, which may depend on a variety of convoluted factors. Therefore, this section performs sensitivity analysis on a range of possible discount factors for the CSO models. The consolidation level, calculated as the percentage of units that are batched (i.e., sourced and shipped by the same location-carrier pair) in all orders, is introduced to measure the effectiveness of the methods in identifying item batching opportunities for cost discounts within each order.
Figure 9 illustrates the performance of C-SAA, C-RO-B, and C-RO-D under a series of different discount factors. As shown in Figure 9(a), C-SAA still consistently exhibits the lowest objective value across all discount factors, with C-RO-B comes next. As expected, all three methods exhibit lower objective values due to lower fulfillment costs as the discount factor increases (see Figure 9(b)). This aligns with the intuition that higher discounts encourage more cost-efficient fulfillment strategies.
The decrease in both C-RO methods can be explained by their higher consolidation levels as the discount factors increases. C-RO methods have higher fulfillment costs, so they benefit more from discounts and capitalize on this opportunity.
The trend in consolidation level for C-SAA is less straightforward. While it peaks at a discount factor of 0.25, the consolidation level remains lower than the baseline throughout. This suggests that, despite some consolidation opportunities, the C-SAA model does not exploit them to the same extent as the C-RO models. One possible explanation is that the benefits of consolidation in C-SAA might not be sufficient to outweigh the penalties associated with deviations. The peak at a discount factor of 0.25 could be an anomaly or indicative of specific conditions where consolidation was briefly more favorable, but overall, the model prioritizes minimizing deviation penalties over achieving higher consolidation.
(a) Performance in Realized Objective Value.
(b) Performance in Fulfillment Costs.
(c) Performance in Consolidation Level.
Figure 9: Performance of CSO Models Across Different Discount Factors.
6 Conclusion
This paper presented a data-driven contextual stochastic optimization framework, which was applied to an industrial omnichannel order fulfillment problem within the context of multi-courier logistics. Leveraging contextual optimization effectively addresses the inherent complexities of modern retail logistics, providing innovative solutions tailored to the dynamic e-commerce landscape. In particular, the proposed framework is capable of handling the availabilities of observational data only.
To achieve these results, the paper uses probabilistic multi-class classification and quantile regression models to estimate the conditional distribution of delivery time deviations. These machine learning models are then seamlessly integrated with SAA-based risk-neutral optimization models and robust optimization models under the proposed CSO framework. The paper evaluated the proposed methodology on a real dataset from one of the largest e-commerce company. The results, which significantly improve the baseline solution, showcase the ability of the CSO framework to improve service quality, customer satisfaction, and operational efficiency. They offer valuable insights and tools for practitioners aiming to optimize their order fulfillment strategies in the evolving e-commerce environment.
There are several potential directions for future work. The contextual optimization models could be extended to incorporate long-term inventory planning through multi-period order fulfillment optimization. This would shift the approach from a rather myopic perspective to a more comprehensive strategy that accounts for future inventory needs. The distributional forecasts of deviations could be improved by increasing the dataset to include factors such as seasonal trends, weather conditions, and carrier tracking details. Furthermore, integrating downstream optimization costs into the training of distributional forecasting models can be an interesting extension (Qi et al. 2021, Elmachtoub and Grigas 2022). However, it is crucial to address the observational nature of the dataset and the varying dimensions of different optimization problem instances. Another promising direction is to extend the current contextual optimization framework to an end-to-end optimization approach (Kotary et al. 2021) while adhering to all operational constraints .
This paper is relevant to the broader literature on uncertainty quantification, where conformal predictions offer a promising direction. Conformal predictions provide coverage guarantees for prediction intervals and could be particularly effective when combined with quantile regression (Romano et al. 2019). Moreover, adopting conformally calibrated predictive distributions (Manokhin 2017) and improving probability calibration methods could enhance model performance. The paper hints at this potential by adopting probability calibration in multi-class classification. The Venn-Abers calibrator was experimented, but it was too computationally intensive to be fitted for the dataset used in the paper.
[…]
This research was partly supported by the NSF AI Institute for Advances in Optimization (Award 2112533).